# In-Context Learning Dynamics with Random Binary Sequences

### v1

@@ Abstract
Large language models (LLMs) trained on huge corpora of text datasets demonstrate complex, emergent capabilities, achieving state-of-the-art performance on tasks they were not explicitly trained for. The precise nature of LLM capabilities is often unclear, and different prompts can elicit different capabilities through in-context learning. We propose a Cognitive Interpretability framework that enables us to analyze in-context learning dynamics to understand latent concepts in LLMs underlying behavioral patterns. This provides a more nuanced understanding than success-or-failure evaluation benchmarks, but does not require observing internal activations as a mechanistic interpretation of circuits would require. Inspired by the cognitive science of human randomness perception, we use random binary sequences as context and study dynamics of in-context learning by manipulating properties of context data, such as sequence length. In the latest GPT-3.5+ models, we find emergent abilities to generate pseudo-random numbers and learn basic formal languages, with striking in-context learning dynamics where model outputs transition sharply from pseudo-random behaviors to deterministic repetition.
@@ 1 Introduction
Large language models (LLMs) achieve high performance on a wide variety of tasks they were not explicitly trained for, demonstrating complex, emergent capabilities (Brown et al., 2020; Radford et al., 2019; Wei et al., 2022a; 2023; b; Lu et al., 2023; Bommasani et al., 2021; Wei et al., 2021; Kojima et al., 2022; Pan et al., 2023; Pallagani et al., 2023; Li et al., 2022). In-context learning (ICL) describes how task-specific patterns of behavior are elicited in LLMs by different prompts (or contexts(Brown et al., 2020; Wei et al., 2022b; Dong et al., 2022; Zhou et al., 2022; Wang et al., 2023; Shin et al., 2022; Min et al., 2022; Si et al., 2023; Xie et al., 2022; Zhang et al., 2023b). Although no weight updates occur in ICL, different input contexts can activate, or re-weight, different latent algorithms in an LLM, analogous to how traditional learning methods such as gradient descent use training data to re-weight model parameters to learn representations (Goldblum et al., 2023; Dai et al., 2023; Ahn et al., 2023; Von Oswald et al., 2023; Akyürek et al., 2022; Li et al., 2023a). Two seemingly equivalent prompts can, however, evoke very different behaviors in LLMs (Si et al., 2023). ICL makes interpreting what representations in an LLM are being evoked by a given input more challenging than interpreting representations in task-specific language models, where the space of capabilities is more restricted and well-defined. Our central motivation here is to interpret emergent capabilities and latent concepts underlying complex behaviors in LLMs by analyzing in-context learning behavioral dynamics, without directly observing hidden unit activations or re-training models on varied datasets.
@@
Inspired by computational approaches to human cognition (Tenenbaum, 1998; Griffiths & Tenenbaum, 2003; Piantadosi et al., 2012; Spivey et al., 2009), we model and interpret latent concepts evoked in LLMs by different contexts, without observing or probing model internals. This approach, which we call Cognitive Interpretability (see Fig. 1), is a middle ground between shallow test-set evaluation benchmarks on one hand (Srivastava et al., 2022; Saparov & He, 2022; Min et al., 2022; Lanham et al., 2023; Ganguli et al., 2023; Bowman et al., 2022; Kadavath et al., 2022; Perez et al., 2022; Prystawski & Goodman, 2023; Turpin et al., 2023) and mechanistic neuron- and circuit-level understanding of pre-trained models’ capabilities on the other (Nanda et al., 2023; Goh et al., 2021; Geva et al., 2020; Belinkov, 2022; Li et al., 2022; Wang et al., 2022; Chughtai et al., 2023; Gurnee et al., 2023; Foote et al., 2023; Lubana et al., 2023; Lieberum et al., 2023; Barak et al., 2022; Liu et al., 2022b; Zhong et al., 2023). The reliance on benchmarks loosely parallels early behaviorist psychology, when theories of human and animal learning only assumed stimulus-response associations, without positing theories about mental processes or neural substrates (Gardner, 1987). Circuit-level mechanistic interpretations of neural networks parallel neurobiology, which offers physical models of information processing in biological neurons, but typically does not account for the structure of high-level behavior. Cognitive interpretability, like cognitive science, is aimed at predicting behavioral outcomes over a potentially infinite space of possible tasks. It defines high-level specifications of behaviors performed by LLMs, which, we argue, should be the first step before mechanistic understanding of circuits existent in a model is pursued.
@@
Here, we focus on cognitive interpretability of LLMs in the domain of random sequences of binary values. Random binary sequences are a minimal domain that have been studied extensively in statistics, formal language theory, and algorithmic information theory (Sipser, 1996; Chaitin, 1966; Li et al., 1997). We can use this domain to systematically test few-shot learning as a function of context length $|x|$ by testing different input sequences $x$. We can also test zero-shot learning by having models generate sequences with no context ($|x|=0$), without relying on alternate prompt formats such as chain-of-thought reasoning (Kojima et al., 2022; Wei et al., 2022b). Moreover, language generation trajectories over binary sequences can also be analyzed and visualized much more easily than typical user-chatbot interaction trajectories (Zhang et al., 2023a; Berglund et al., 2023), since the token-by-token branching factor is only two. Random binary sequences have also been a target domain in cognitive science (specifically, subjective randomness), where researchers have studied the mechanisms and concepts that underlie how people generate random binary sequences or evaluate the randomness of given sequences (Falk & Konold, 1997; Tversky & Gilovich, 1989; Griffiths et al., 2018).
Figure 1: Cognitive Interpretability in the context of prior work Cognitive Interpretability studies in-context learning dynamics in LLMs, positing theories of the latent concepts enabling behavioral capabilities. It is a middle ground between behavioral benchmarks, which treat models as black boxes and evaluate hit-or-miss accuracy, and mechanistic interpretability, which studies toy models and training loss dynamics, analogous to how cognitive science is a middle ground between behaviorist psychology and neurobiology in the study of human intelligence.
Computational cognitive scientists have related algorithmic information theory to human cognition, where mental concepts are viewed as programs, and cognitive hypothesis search over concepts is viewed as Bayesian inference (or Bayesian program learning) (Chater & Vitányi, 2003; Dehaene et al., 2022; Piantadosi et al., 2012; Goodman et al., 2008; Ullman et al., 2012; Ullman & Tenenbaum, 2020). In this vein, Griffiths & Tenenbaum (2003) and Griffiths et al. (2018) model subjective randomness in human cognition as probabilistic program induction, where a person must search over a space of non-random programs in order to answer the question, “was this sequence generated by a random process?” We argue that ICL can similarly be seen as under-specified program induction, where there is no single “correct” answer; instead, an LLM should appropriately re-weight latent algorithms. The domain of random sequences reflects this framing, in contrast to other behavioral evaluation methodologies, in that there is no correct answer to a random number generation or judgment task (Fig. 3). If the correct behavior is to match a target random process, then the right way to respond to a prompt Generate N flips from a fair coin is at best a uniform distribution over the tokens Heads and Tails, instead of a specific ground truth token sequence. If instead the correct behavior is to match how humans generate random sequences, the LLM must represent a more complex algorithm that matches the algorithm presumably being implemented by people. Overall, we make the following contributions in this work:
  • Evaluating ICL as Bayesian model selection in a practical setting. We find evidence of in-context learning operating as Bayesian model selection in LLMs in the wild, compared to prior work that trains smaller toy neural networks from scratch.
  • Subjective Randomness as a domain for studying ICL dynamics. We introduce a new domain, subjective randomness, which is both rich enough to be theoretically interesting, but simple enough that complex behavioral patterns are easily analyzed.
  • Sharp transitions in abilities to generate pseudo random numbers. We systematically analyze in-context learning dynamics in LLMs without observing or updating model weights, demonstrating sharp phase-changes in model behavior. This is a minimal, interpretable example of how the largest and most heavily fine-tuned LLMs can suddenly shift from one pattern of behavior to another during text generation.
2 Preliminaries: Bayesian Inference and In-Context Learning
@@
Figure 2: Overview of our modeling framework. Given a pre-trained Large Language Model, we systematically vary input context prompts, in this case $x=(001)^{n}$ with $n=\{2,4\}$ (001001, $001$). The LM outputs $y$ (red) vary significantly as a function of $x$ (blue), based on some unknown latent concept space embedded in the LLM. We model a subset of the algorithms embedded in the LLM’s latent concept space $\mathcal{H}$ that are invoked by ICL (red + shaded box), to predict the LLM outputs $y$. With very little context ($x=\texttt{0}$), GPT-3.5+ generate subjectively random sequences, whereas with adequate context matching a simple formal language ($x=\texttt{001001001001}$), behavior follows the same deterministic pattern.
A multi-layer neural networks may possess multiple distributed circuits implementing computational primitives, such as basic mathematical operations like addition and sequence copying (Olsson et al., 2022; Nanda et al., 2023; Hanna et al., 2023; Zhong et al., 2023; Chan et al., 2022; Bietti et al., 2023). In state of the art LLMs with hundreds of billions (even trillions) of parameters, there may be sub-networks implementing various computations, and a wide variety of emergent behaviors corresponding to those computations. A similar situation occurs in research focused on understanding the human brain, where sub-networks of neuronal cells have been shown to localize specific capabilities. In cognitive science, researchers study such aspects of cognition without fully understanding the underlying neural circuitry, often modeling behavior without observing brain activity. Analogously, we seek to understand the structure of behaviors in large language models in the wild, without a full mechanistic understanding of their circuitry.
@@
Bayesian inference is a key methodological tool of cognitive modeling, and recent work has framed in-context learning as Bayesian inference over models (Xie et al., 2022; Li et al., 2023b; Hahn & Goyal, 2023). Specifically, the posterior predictive distribution $p(y|x)$ in these works describes how an LLM produces output tokens $y$, given the context tokens, or prompt, $x$. The key assumption is that a context $x$ will activate latent concepts $c$ within a model according to their posterior probability $p(c|x)$, which the model marginalizes over to produce the next token $y$ by sampling from the posterior predictive distribution: $p(y|x)=\int_{c\in\mathcal{C}}p(y|c)\ p(c|x)$. This Bayesian model selection takes place in network activation dynamics, without changing model weights.
@@
Following the framing above, in our experiments we assume a hypothesis space $\mathcal{H}$ that approximates the latent space of LLM concepts used when predicting the next token, i.e.,
@@
where $\sum_{h}$ can be changed to $\max_{h}$ to represent deterministic greedy decoding with an LLM temperature parameter of 0. If we assume that an LLM operates over next-token prediction only, then each generated token $y_{t-1}$ will be appended to the context when the following token is sampled, i.e. $y_{t}\sim p(y_{t}|x,y_{t-1})$. We specifically focus on Bernoulli processes, regular languages, Markov chains, and a simple memory-constrained probabilistic model as candidates for the hypothesis space $\mathcal{H}$ for estimating LLM concepts in random binary sequences. These candidates together capture a variety of behavioral patterns with varying complexity: Bernoulli processes serve as a baseline for a true random coin flip, where all elements in a sequence are independent and identically distributed; Markov chains are a minimal model capable of representing serial correlation patterns; the Gambler’s Fallacy bias in cognitive science (described below) emerges from a simple memory-limited model that draws the sequence towards a specified probability dependent based on previous samples; and Regular languages are well studied in theoretical computer science (Sipser, 1996; Chomsky, 1956), representing a simple class of non-random programs with known mechanisms of finite state automata (or, equivalently, finite state Markov processes). We use a subset of regular languages $(x)^{n}$, where $(x)$ is a short sequence of values, e.g., $(\texttt{010})^{n}$, where 00 maps to Heads and $1$ to Tails.
Learning Dynamics in Model Selection
@@
Cognitive scientists have used Bayesian predictive and posterior distributions to model learning dynamics in human adults and children, as well as in non-human animals (Tenenbaum, 1998; Goodman et al., 2008; Ullman & Tenenbaum, 2020). A discrete hypothesis space in a probabilistic model can give clear and meaningful explanations to learning patterns analogous to mode collapse and phase transitions in deep learning. When behavior suddenly shifts from one pattern, or mode, of behavior to another, this can be understood as one hypothesis coming to dominate the posterior $p(h|x)$ as $x$ grows in scale (Ullman et al., 2012). For example, children between the ages of 3.5—5 years old learning to count undergo a dramatic conceptual shift from knowing the meanings of only a few number words (“one”, “two”) to a full inductive understanding of counting, which can be modeled as Bayesian model selection with a simplicity prior over models (Piantadosi et al., 2012). Such cognitive analysis of behavioral shifts as generated from a shift in posterior probability between one hypothesis to another parallels recent work on learning dynamics in training and prompting LLMs. E.g., sharp phase changes have been observed when training neural networks, corresponding to the formation of identifiable mechanisms such as modular addition circuits or induction heads (Nanda et al., 2023; Olsson et al., 2022; Zhong et al., 2023; Bai et al., 2023). ICL research has explored how few-shot prompting can significantly boost LLM performance in various domains, and how the particular exemplars provided in context determine its overall effectiveness (Brown et al., 2020; Zhao et al., 2020; 2021; Min et al., 2022; Lu et al., 2021; Zhou et al., 2022; Turpin et al., 2023). Work on chain-of-thought reasoning in LLMs demonstrates how a few exemplars of detailed solutions or even a simple prompt like “let’s think this through step-by-step” can dramatically impact model performance (Wei et al., 2022b; Kojima et al., 2022; Srivastava et al., 2022). For further discussion of related work, see Appendix A.

### v3

@@ Abstract
Large language models (LLMs) trained on huge text datasets demonstrate intriguing capabilities, achieving state-of-the-art performance on tasks they were not explicitly trained for. The precise nature of LLM capabilities is often mysterious, and different prompts can elicit different capabilities through in-context learning. We propose a framework that enables us to analyze in-context learning dynamics to understand latent concepts underlying LLMs’ behavioral patterns. This provides a more nuanced understanding than success-or-failure evaluation benchmarks, but does not require observing internal activations as a mechanistic interpretation of circuits would. Inspired by the cognitive science of human randomness perception, we use random binary sequences as context and study dynamics of in-context learning by manipulating properties of context data, such as sequence length. In the latest GPT-3.5+ models, we find emergent abilities to generate seemingly random numbers and learn basic formal languages, with striking in-context learning dynamics where model outputs transition sharply from seemingly random behaviors to deterministic repetition.
**footnotetext: Correspondence to: ebigelow@g.harvard.edufootnotetext: Equal contribution
@@ 1 Introduction
Large language models (LLMs) achieve high performance on a wide variety of tasks they were not explicitly trained for, demonstrating complex, emergent capabilities (Brown et al., 2020; Radford et al., 2019; Wei et al., 2022a; 2023; b; Lu et al., 2023; Bommasani et al., 2021; Wei et al., 2021; Kojima et al., 2022; Pan et al., 2023; Pallagani et al., 2023; Li et al., 2022). In-context learning (ICL) describes how task-specific patterns of behavior are elicited in LLMs by different prompts (or contexts(Brown et al., 2020; Wei et al., 2022b; Dong et al., 2022; Zhou et al., 2022; Wang et al., 2023; Shin et al., 2022; Min et al., 2022; Si et al., 2023; Xie et al., 2022; Zhang et al., 2023b). Although no weight updates occur in ICL, different input contexts can activate, or re-weight, different latent algorithms in an LLM, analogous to updating model parameters to learn representations in traditional gradient descent learning methods (Goldblum et al., 2023; Dai et al., 2023; Ahn et al., 2023; Von Oswald et al., 2023; Akyürek et al., 2022; Li et al., 2023a). However, two seemingly equivalent prompts can evoke very different behaviors in LLMs (Si et al., 2023). ICL makes interpreting what representations in an LLM are being evoked by a given input more challenging than with task-specific language models, where the space of capabilities is more restricted and well-defined.
@@
Our central motivation is to interpret latent concepts underlying complex behaviors in LLMs by analyzing in-context learning behavioral dynamics, without directly observing hidden unit activations or re-training models on varied datasets. Inspired by computational approaches to human cognition (Tenenbaum, 1998; Griffiths & Tenenbaum, 2003; Piantadosi et al., 2012; Spivey et al., 2009), we model and interpret latent concepts evoked in LLMs by different contexts without observing or probing model internals. This approach, which we call Cognitive Interpretability, is a middle ground between shallow test-set evaluation benchmarks on one hand (Srivastava et al., 2022; Saparov & He, 2022; Min et al., 2022; Lanham et al., 2023; Ganguli et al., 2023; Bowman et al., 2022; Kadavath et al., 2022; Perez et al., 2022; Prystawski & Goodman, 2023; Turpin et al., 2023) and mechanistic neuron- and circuit-level understanding of pre-trained model capabilities on the other (Nanda et al., 2023; Goh et al., 2021; Geva et al., 2020; Belinkov, 2022; Li et al., 2022; Wang et al., 2022; Chughtai et al., 2023; Gurnee et al., 2023; Foote et al., 2023; Lubana et al., 2023; Lieberum et al., 2023; Barak et al., 2022; Liu et al., 2022b; Zhong et al., 2023) (also see App. C). Unlike related work of feature interpretability in vision neural networks, we analyze ICL and text generation dynamics over output sequences, and unlike mechanistic work studying ICL, we observe learning dynamics in black box input-output behavior of state-of-the-art LLMs.
@@
Specifically, we focus on interpreting latent concepts in LLMs in the domain of random sequences of binary values (Fig. 1). Random binary sequences are a minimal domain that has been studied extensively in statistics, formal language theory, and algorithmic information theory (Sipser, 1996; Chaitin, 1966; Li et al., 1997). We use this domain to systematically test few-shot learning as a function of context length $|x|$ across different binary token sequences $x$. Moreover, language generation trajectories over binary sequences can also be analyzed and visualized more easily than typical user-chatbot interaction trajectories (Zhang et al., 2023a; Berglund et al., 2023) since the token-by-token branching factor is only two. Random binary sequences have also been a target domain in cognitive science (specifically, subjective randomness), where researchers have studied the mechanisms and concepts that underlie how people generate random binary sequences or evaluate the randomness of a sequence (Falk & Konold, 1997; Tversky & Gilovich, 1989; Griffiths et al., 2018).
@@
Figure 1: Overview of our modeling framework. Given a pre-trained Large Language Model, we systematically vary input context prompts, in this case $x=(001)^{n}$ with $n=\{2,4\}$ (001001, $001$). LLM outputs $y$ (Bottom) significantly as a function of $x$ (Top), based on some unknown latent concept space embedded in the LLM. We model a subset of the algorithms embedded in the LLM’s latent concept space $\mathcal{H}$ that are invoked by ICL (Middle), to predict the LLM outputs $y$. With very little context (Left), GPT-3.5+ generates subjectively random sequences, whereas with enough context matching a simple formal language ($x=\texttt{001001001001}=(\texttt{001})^{4}$), it begins deterministically repeating patterns in $x$.
Similar to computational cognitive theories of human concept learning (Chater & Vitányi, 2003; Dehaene et al., 2022; Piantadosi et al., 2012; Goodman et al., 2008; Ullman et al., 2012; Ullman & Tenenbaum, 2020), we argue that ICL can be interpreted as under-specified program induction, where there is no single “correct” answer; instead, an LLM should appropriately re-weight latent algorithms. The domain of random sequences reflects this framing, in contrast to other behavioral evaluation methodologies, in that there is no correct answer to a random number generation or judgment task (Fig. 3). If the correct behavior is to match a target random process, then the best way to respond to the prompt Generate N flips from a fair coin is a uniform distribution over the tokens Heads and Tails, instead of a specific ground truth token sequence. If instead the correct behavior is to match how humans generate random sequences, the LLM must represent a more complex algorithm that matches the algorithm used by people. Overall, we make the following contributions in this work.
  • Evaluating ICL as Bayesian model selection in a practical setting (Sec. 2). We find evidence of in-context learning operating as Bayesian model selection in LLMs in the wild, compared to prior work that trains smaller toy neural networks from scratch.
  • Subjective Randomness as a domain for studying ICL dynamics (Sec. 3). We introduce a new domain, subjective randomness, which is rich enough to be theoretically interesting, but simple enough to enable analysis of complex behavioral patterns.
  • Sharp transitions in abilities to in-context learning dynamics (Fig. 4,  5). We systematically analyze in-context learning dynamics in LLMs without observing or updating model weights, demonstrating sharp phase-changes in model behavior. This is a minimal, interpretable example of how the largest and most heavily fine-tuned LLMs can suddenly shift from one pattern of behavior to another during text generation, and supports theories of ICL as model selection.
2 In-Context Learning as Model Selection
@@
Bayesian inference is a key methodological tool of cognitive modeling, and recent work has framed in-context learning as Bayesian inference over models (Xie et al., 2022; Li et al., 2023b; Hahn & Goyal, 2023). The output tokens $y$ an LLM produces given input tokens $x$ can be modeled as a posterior predictive distribution $p(y|x)$ that marginalizes over latent concepts $c$ in the LLM: $p(y|x)=\int_{c\in\mathcal{C}}p(y|c)\ p(c|x)$ In other words, a context $x$ will activate latent concepts $c$ in an LLM according to their posterior probability $p(c|x)$, which the model marginalizes over to produce the next token $y$ by sampling from the posterior predictive distribution. This inference process takes place in network activation dynamics, without changing model weights.
Figure 2: Sharp changes in the predictive distribution $p(y|x)$ suggest model selection (Left) Here we show an illustrative example of a simple Bayesian model with two hypotheses in $\mathcal{H}$: a random Bernoulli hypothesis, and a deterministic Non-Random hypothesis for the concept $(\texttt{01})^{n}$. When hypothesis posteriors $p(h|x)$ cross as a function of additional data, i.e. larger $|x|$, the predictive distribution $p(y|x)$ for model selection suddenly transitions from one pattern of behavior to another (Right). With model averaging, instead there is a steady, gradual change in $p(y|x)$ (see App. D for further explanation). We find analogous S-shaped curves in ICL with LLM predictive distributions $p(y|x)$ in GPT-3.5+, suggesting model selection rather than model averaging (Fig. 4,  5). These phase changes are not predicted by theories of ICL as linear regression (Akyürek et al., 2022) or few-shot learning (Brown et al., 2020; Wei et al., 2022b), where loss steadily decreases with more context examples in $x$.
@@
In our experiments, we assume a hypothesis space $\mathcal{H}$ that approximates the latent space of LLM concepts used when predicting the next token, i.e.,
@@
We assume that an LLM operates over next-token prediction only, and so each generated token $y_{t-1}$ will be appended to the context when the following token is sampled, i.e. $x_{t}=(x_{0,\ldots,t-1},y_{t-1})$. With this framework, we can understand ICL dynamics through the lens of hypothesis search. An LLM might not compute the sum in the above equation by taking a weighted average of latent capabilities. An LLM might perform Bayesian model selection, instead of model averaging, where for any given $x$, only a single latent concept $h$ approximates the posterior predictive distribution: $h^{*}=\operatorname*{argmax}_{h}p(h|x),\ \ p(y|x)=p(y|h^{*})$. We provide a visual demonstration of how Bayesian model selection leads to sharp changes in the predictive distribution $p(y|x)$ with a toy model in Fig. 2.
@@
Cognitive scientists have used Bayesian predictive and posterior distributions to model learning dynamics in human adults and children, as well as in non-human animals (Tenenbaum, 1998; Goodman et al., 2008; Ullman & Tenenbaum, 2020). A discrete hypothesis space in a probabilistic model can give clear and meaningful explanations to learning patterns analogous to mode collapse and phase transitions in deep learning. When behavior suddenly shifts from one pattern, or mode, of behavior to another, this can be understood as one hypothesis coming to dominate the posterior $p(h|x)$ as $x$ grows in scale. This theory can explain conceptual shifts observed when children learn intuitive theories of domains such as intuitive physics (Ullman et al., 2012) and learning to count  (Piantadosi et al., 2012). Such cognitive analysis of behavioral shifts as generated from a shift in posterior probability between one hypothesis to another parallels recent work on learning dynamics in training and prompting LLMs. Sharp phase changes have been observed when training neural networks, corresponding to the formation of identifiable mechanisms such as modular addition circuits or induction heads (Nanda et al., 2023; Olsson et al., 2022; Zhong et al., 2023; Bai et al., 2023), and can be interpreted as Bayesian model selection (Hoogland et al., ). Also see App. AC.
To our knowledge, we provide the first demonstration of such sharp phase changes in ICL, particularly with state-of-the art blackbox LLMs. The minimal model used by Xie et al. (2022), as well as other few-shot learning work (Brown et al., 2020), show linearly or logarithmically decreasing loss curves in ICL, rather than S-shaped curves with stable plateaus as we find. Michaud et al. (2023) propose that neural network knowledge and skills may be quantized into discrete chunks, and so training might consist of many such phase changes which disappear into a smooth curve when test loss is averaged over many samples. We propose that In-Context Learning may also consist of sudden phase changes, as well as other complex learning dynamics, which may be explained as search over a discrete latent concept space. Such phase changes may occur in typical few-shot learning domains (Wei et al., 2022b; Brown et al., 2020), where smooth, continuously decreasing loss curves might hide complex ICL dynamics corresponding to shifts in the latent algorithms driving an LLMs behavior.